0 5 7 0 0 emmme lu 643? 1317N Wh A B C D
Answer #1

7.35 (\frac{v}{v}) % of Sucrose solution means 7.35 ml of Sucrose is dissolved in 100 ml solution.

concentration of sucrose ( \frac{v}{v} %) = \frac{volume \: of \: sucrose \; (mL) \times 100 }{volume\; of \; solution (mL)}

when, volume of the solution = 2.00 L = 2000 mL

Concentration of Sucrose = 7.35 %

then, volume of Sucrose = (volume of Solution \times concentration of Sucrose)/100

= 2000\times7.35/100 = 147 mL

Hence right option is B. 147 mL